# Unsupervised Video Domain Adaptation with Masked Pre-Training and Collaborative Self-Training

### v3

@@ 2 Related Work
Video Unsupervised Domain Adaptation (VUDA). Techniques for video unsupervised domain adaptation typically fall within a few common categories [59]. Adversarial methods attempt to align representations between the source and target domains by enforcing a domain confusion loss using a separate domain discriminator network, where the main network must learn to fool the discriminator. Some techniques which were developed for images, such as DANN [18] or MK-MMD [34], can be applied directly to video architectures. TAN [8] is a video-specific approach that uses separate adversaries for the spatial and temporal dimensions. These approaches can result in unstable training due to competing objectives.
Another class of methods use contrastive learning and exploit the intrinsic structure of video. CoMix [45] uses contrastive learning with pseudolabeling and video-based augmentations such as sampling at varying frame rates or mixing the backgrounds of source videos and target videos. COA [52] optimizes 6 different losses simultaneously, including contrastive losses at the clip- and video-level, along with supervised contrastive learning across domains. UDAVT [10] adapts vision transformers using the information bottleneck principle by having separate projectors for source and target features and estimating a cross correlation matrix between features whose labels and pseudolabels match.
@@ 3.2 Self-Supervised Initialization
In contrast to some other video domain adaptation works [10, 65] that use network weights from Kinetics-400 supervised pre-training, we initialize our network from self-supervised pre-training. As contended in [43] for UDA with images, we argue that supervised pre-training can potentially complicate the study of VUDA techniques. When the pre-trained network has been trained to classify categories present in the DA dataset, some DA techniques could perform well simply by preserving the capabilities of the pre-trained model despite not generalizing well to DA tasks that do not have category overlap with the pre-training dataset. Moreover, the UDA problem formulation implies that labeled instances of relevant classes are only available from the source domain. This condition does not hold for the Daily-DA and Sports-DA benchmarks that we use for evaluation, as 6 out of their 8 classes also appear in the Kinetics-400 dataset. Further exacerbating the issue with supervised Kinetics pre-training is the fact that Kinetics serves as a target domain in both benchmarks. Thus, we initialize the network weights from UMT [30] single-modality (i.e., video only) self-supervised pre-training on Kinetics-710 [29]. Please refer to Sec. 10 in the Appendix for select results using a UMT pre-trained network with additional supervised fine-tuning on Kinetics-400, which unsurprisingly achieves better baseline performance.
@@ 5.1 Datasets
We evaluate our approach on three video domain adaptation benchmarks: Daily-DA, Sports-DA, UCFHMDB.
@@
ZS |
CLIP (ViT-B/16) [44] | 36.5 | 36.5 | 36.5 | 60.0 | 60.0 | 60.0 | 48.5 | 48.5 | 48.5 | 68.1 | 68.1 | 68.1 | 53.3 |
SFVUDA |
ATCoN [62] | 17.9 | 27.2 | 17.2 | 26.7 | 47.3 | 48.2 | 30.7 | 17.2 | 32.5 | 57.7 | 48.5 | 31.0 | 33.5 |
@@
ZS |
CLIP (ViT-B/16) [44] | 85.0 | 85.0 | 93.3 | 93.3 | 91.3 | 91.3 | 89.9 |
SFVUDA |
ATCoN [62] | 47.9 | 69.7 | 90.6 | 93.6 | 65.2 | 76.0 | 73.8 |
@@ 5.3 Baselines
We compare results using UNITE with previously developed techniques in UDA, as well as in the closely related setting of SFUDA as reported by [67]. For Daily-DA and Sports-DA we compare against DANN [18], MK-MMD [34] and TAN [8] in the UDA setting, and ATCoN [62], EXTERN [63] and DALL-V [67] for SFUDA. For UCFHMDB, we additionally report accuracies from COA [52] and UDAVT [10].
@@ 5.4 Benchmark Results
Tables 1, 2 and 3 show the results of our evaluation of UNITE on Daily-DA, Sports-DA and UCFHMDB. We include accuracies after Stage 2 of the pipeline (denoted as UNITE w/o CST) as well as after applying the full UNITE process. For reference, we include accuracies using our UMT pre-trained network with source domain supervised fine-tuning (Source Only) and with target domain supervised fine-tuning (Target Only). We also include the source only and target only accuracies (denoted in the tables as LB and UB) of DALL-V [67] (a CLIP-based ResNet-50 model) for added context. We additionally display the zero-shot classification accuracy of the CLIP teacher model that we utilize during Stage 1 and Stage 3 of UNITE.
@@
On UCFHMDB we see state-of-the-art performance on UH, while UDAVT outperforms on HU. Nevertheless, UNITE still achieves strong overall performance on the benchmark.
@@
ZS |
CLIP (ViT-B/16) [44] | 88.8 | 91.7 | 90.3 |
SFVUDA |
ATCoN [62] | 85.3 | 79.7 | 82.5 |
@@ 6 Additional Analysis & Discussion
Student exceeds the teacher. We see in the benchmark results that the video student model exceeds the zero-shot accuracy of the CLIP teacher in virtually all domain shifts. This suggests that UNITE successfully leverages the spatial modeling capabilities of CLIP to train an even more powerful spatiotemporal model for the action recognition task. We can see from Figure 4 that the teacher model and the student model prior to collaborative self-training have differing capabilities. For example, the image teacher is stronger at recognizing actions from the jump class than the Stage 2 student. On the other hand, the video student is significantly better than the teacher at recognizing wave at the start of CST. In both cases, we see that the student model, through collaborative self-training, comes to exceeds both the Stage 2 student and the CLIP image teacher. With the MatchOrConf pseudolabeling scheme, the two models are able to work together to achieve higher accuracy in the target domain. Accuracies on the pick and run classes decline after CST, as self-training tends to reinforce the accuracy of well-performing classes, sometimes at the expense of more difficult ones that are infrequently assigned pseudolabels. Modifying CST to account for the learning status of each class, like in [68], may help remedy this issue.
@@
Comparison of pseudolabeling strategies. In Table 7, we explore the use of alternative pseudolabeling strategies to the MatchOrConf [69] scheme that we employ in UNITE. The first 3 rows of the table assess variations of the masked consistency pseudolabeling scheme proposed in PACMAC [43], where target samples are selected if there is agreement between predictions of masked views. This criterion can also be combined with a constraint on the confidence exceeding threshold . While we find the combination of these two constraints (ConsOrConf) to be effective for self-training, we find that MatchOrConf results in more consistent target accuracy improvements across domain shifts. We also consider combining MatchOrConf with a masked consistency constraint as a way to reduce the rate of errors where the student and teacher models agree on an incorrect prediction. This scheme, denoted as MatchAndConsOrConf adds an additional constraint to the matching criterion by also requiring agreement between the student model predictions of masked views. We find that this approach underperforms compared to the vanilla MatchOrConf scheme due to low utilization of target domain samples.
@@ 7 Conclusions
Acknowledgements. This research was sponsored by the Army Research Laboratory and was accomplished under Cooperative Agreement Number W911NF-21-2-0211. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ 11 Zero-Shot Classification with CLIP
| Daily-DA | Sports-DA | UCFHMDB |

### v4

@@ 2 Related Work
Video Unsupervised Domain Adaptation (VUDA). Techniques for video unsupervised domain adaptation typically fall within a few common categories [59]. Adversarial methods attempt to align representations between the source and target domains by enforcing a domain confusion loss using a separate domain discriminator network, where the main network must learn to fool the discriminator. Some techniques which were developed for images, such as DANN [18] or MK-MMD [34], can be applied directly to video architectures. TA3N [8] is a video-specific approach that uses separate adversaries for the spatial and temporal dimensions. These approaches can result in unstable training due to competing objectives.
Another class of methods use contrastive learning and exploit the intrinsic structure of video. CoMix [45] uses contrastive learning with pseudolabeling and video-based augmentations such as sampling at varying frame rates or mixing the backgrounds of source videos and target videos. CO2A [52] optimizes 6 different losses simultaneously, including contrastive losses at the clip- and video-level, along with supervised contrastive learning across domains. UDAVT [10] adapts vision transformers using the information bottleneck principle by having separate projectors for source and target features and estimating a cross correlation matrix between features whose labels and pseudolabels match.
@@ 3.2 Self-Supervised Initialization
In contrast to some other video domain adaptation works [10, 65] that use network weights from Kinetics-400 supervised pre-training, we initialize our network from self-supervised pre-training. As contended in [43] for UDA with images, we argue that supervised pre-training can potentially complicate the study of VUDA techniques. When the pre-trained network has been trained to classify categories present in the DA dataset, some DA techniques could perform well simply by preserving the capabilities of the pre-trained model despite not generalizing well to DA tasks that do not have category overlap with the pre-training dataset. Moreover, the UDA problem formulation implies that labeled instances of relevant classes are only available from the source domain. This condition does not hold for the Daily-DA and Sports-DA benchmarks that we use in our evaluations, as 6 out of their 8 classes also appear in the Kinetics-400 dataset. Further exacerbating the issue with supervised Kinetics pre-training is the fact that Kinetics serves as a target domain in both benchmarks. Thus, we initialize the network weights from UMT [30] single-modality (i.e., video only) self-supervised pre-training on Kinetics-710 [29]. Please refer to Sec. 10 in the Appendix for select results using a UMT pre-trained network with additional supervised fine-tuning on Kinetics-400, which unsurprisingly achieves better baseline performance.
@@ 5.1 Datasets
We evaluate our approach on three video domain adaptation benchmarks: Daily-DA, Sports-DA, UCFHMDBfull.
@@
| ZS | CLIP (ViT-B/16) [44] | 36.5 | 36.5 | 36.5 | 60.0 | 60.0 | 60.0 | 48.5 | 48.5 | 48.5 | 68.1 | 68.1 | 68.1 | 53.3 |
| SFVUDA | ATCoN [62] | 17.9 | 27.2 | 17.2 | 26.7 | 47.3 | 48.2 | 30.7 | 17.2 | 32.5 | 57.7 | 48.5 | 31.0 | 33.5 |
@@
| ZS | CLIP (ViT-B/16) [44] | 85.0 | 85.0 | 93.3 | 93.3 | 91.3 | 91.3 | 89.9 |
| SFVUDA | ATCoN [62] | 47.9 | 69.7 | 90.6 | 93.6 | 65.2 | 76.0 | 73.8 |
@@ 5.3 Baselines
We compare results using UNITE with previously developed techniques in UDA, as well as in the closely related setting of SFUDA as reported by [67]. For Daily-DA and Sports-DA we compare against DANN [18], MK-MMD [34] and TA3N [8] in the UDA setting, and ATCoN [62], EXTERN [63] and DALL-V [67] for SFUDA. For UCFHMDBfull, we additionally report accuracies from CO2A [52] and UDAVT [10].
@@ 5.4 Benchmark Results
Tables 1, 2 and 3 show the results of our evaluation of UNITE on Daily-DA, Sports-DA and UCFHMDBfull. We include accuracies after Stage 2 of the pipeline (denoted as UNITE w/o CST) as well as after applying the full UNITE process. For reference, we include accuracies using our UMT pre-trained network with source domain supervised fine-tuning (Source Only) and with target domain supervised fine-tuning (Target Only). We also include the source only and target only accuracies (denoted in the tables as LB and UB) of DALL-V [67] (a CLIP-based ResNet-50 model) for added context. We additionally display the zero-shot classification accuracy of the CLIP teacher model that we utilize during Stage 1 and Stage 3 of UNITE.
@@
On UCFHMDBfull we see state-of-the-art performance on UH, while UDAVT outperforms on HU. Nevertheless, UNITE still achieves strong overall performance on the benchmark.
@@
| ZS | CLIP (ViT-B/16) [44] | 88.8 | 91.7 | 90.3 |
| SFVUDA | ATCoN [62] | 85.3 | 79.7 | 82.5 |
@@ 6 Additional Analysis & Discussion
Student exceeds the teacher. We see in the benchmark results that the video student model exceeds the zero-shot accuracy of the CLIP teacher in virtually all domain shifts. This suggests that UNITE successfully leverages the spatial modeling capabilities of CLIP to train an even more powerful spatiotemporal model for the action recognition task. We can see from Figure 4 that the teacher model and the student model prior to collaborative self-training have differing capabilities. For example, the image teacher is stronger at recognizing actions from the jump class than the Stage 2 student. On the other hand, the video student is significantly better than the teacher at recognizing wave at the start of CST. In both cases, we see that the student model, through collaborative self-training, comes to exceeds both the Stage 2 student and the CLIP image teacher. With the MatchOrConf pseudolabeling scheme, the two models are able to work together to achieve higher accuracy in the target domain. We see accuracy on the pick and run classes declines after CST, as self-training tends to reinforce the accuracy of well-performing classes, sometimes at the expense of more difficult ones that are infrequently assigned pseudolabels. Modifying CST to account for the learning status of each class, like in [68], may help remedy this issue.
@@
Comparison of pseudolabeling strategies. In Table 7, we explore the use of alternative pseudolabeling strategies to the MatchOrConf [69] scheme employed in UNITE. The first 3 rows of the table assess variations of the masked consistency pseudolabeling scheme proposed in PACMAC [43], where target samples are selected if there is agreement between predictions of masked views. This criterion can also be combined with a constraint on the confidence exceeding threshold . While we find the combination of these two constraints (ConsOrConf) to be effective for self-training, we find that MatchOrConf results in more consistent target accuracy improvements across domain shifts. We also consider combining MatchOrConf with a masked consistency constraint as a way to reduce cases where the student and teacher models agree on an incorrect prediction. This scheme, denoted as MatchAndConsOrConf, adds an additional constraint to the matching criterion by also requiring agreement between the student model predictions of masked views. We find that this approach underperforms compared to the vanilla MatchOrConf scheme due to low utilization of target domain samples. Our results also indicate that using both models together for pseudolableing outperforms using only zero-shot CLIP predictions.
@@ 7 Conclusions
Acknowledgements. This research was sponsored by the Army Research Laboratory under Cooperative Agreement W911NF-21-2-0211. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ 11 Zero-Shot Classification with CLIP
| Daily-DA | Sports-DA | UCFHMDBfull |